Calculate the following for the banana republic given the information provided above
a. What is the consumption function?
b. What is the equation of the AE curve?
c. Calculate equilibrium expenditure
d. Calculate the multiplier.
1
Expert's answer
2019-10-18T09:45:22-0400

a. The consumption function is C = C0 + cYd, where C0 is autonomous consumption, c is marginal propensity to consume, Yd is disposable income.

b. The equation of the AE curve is: AE = Y = C + I + G + NX.

c. The equilibrium expenditure is when AE = Y = C + I + G + NX.

d. Calculate the multiplier formula is: m = 1/(1 - c(1 - t)), where c is marginal propensity to consume, t is marginal tax rate.